OpenStudy (anonymous):

Find the value of Y. What are the measures of the missing two angles? [2(y+5)] (9y+5)

Parth (parthkohli):

What's the sign in between there?

OpenStudy (anonymous):

idk it didnt say i think =

Parth (parthkohli):

Oh I get it. \(\Large \color{Black}{\Rightarrow 2y + 10 = 9y + 5}\) \(\Large \color{Black}{\Rightarrow 7y = 5 }\) \(\Large \color{Black}{\Rightarrow y = {-5 \over 7}}\)

OpenStudy (anonymous):

^^^^ divided righ?

Parth (parthkohli):

And then we have to substitute.

Parth (parthkohli):

Yes exactly.

OpenStudy (anonymous):

0.714..............

Parth (parthkohli):

To find 2(y + 5), just substitute the value of y.

OpenStudy (anonymous):

\[2y+10=9y+5 \]\[y= -5/7\] then i get 0.714 but it dont look rihg

Parth (parthkohli):

Umm

OpenStudy (anonymous):

2y+10=9y+5 7y=5 y= -5/7 then i get 0.714..................... (lots of numbs)